# MOLLM: Multi-Objective Large Language Model for Molecular Design – Optimizing with Experts

### v1

@@ 5.2 Experience Pool
Top1 F | Top10 F | Uniqueness | Validity | Diversity | |
0.0 | 4.187 | 4.152 | 0.937 | 0.915 | 0.556 |
0.1 | 4.175 | 4.163 | 0.935 | 0.917 | 0.548 |
0.3 | 4.154 | 4.124 | 0.961 | 0.903 | 0.544 |
0.5 | 4.168 | 4.144 | 0.978 | 0.898 | 0.554 |
Inspired by ExpeL (Zhao et al., 2023), we incorporate an experience pool into our algorithm to enhance molecular optimization. The experience pool consists of two key components: (1) knowledge gained from generating better and structurally similar molecules, and (2) insights for avoiding suboptimal molecules. These are achieved by summarizing information from the top 10 and bottom 10 molecules, respectively, in each iteration. The worst 10 molecules are selected using a sliding window approach with a stride of 10. Specifically, if in the previous iteration, the worst molecules were extracted from the bottom 10, the next iteration extracts molecules ranked from the bottom 20 to the bottom 10. This mechanism ensures that the experience pool continuously evolves, integrating knowledge from both current and past iterations.
While the concept of experience pools aligns with human decision-making—shifting from general heuristics to more concrete optimization strategies—we observe a performance decline when incorporating experience into our model. As shown in Table 4, where represents the probability of integrating experience into the prompt, the best performance is achieved when , indicating that excluding experience leads to superior optimization and greater molecular diversity. We attribute this phenomenon to the nature of optima distribution in the molecular space. Since local optima tend to be large and widely separated, the experience pool may cause the model to focus excessively on a given local optimum, thereby hindering exploration of alternative high-quality solutions. Consequently, to maximize optimization performance, we temporarily exclude the experience pool from our experiments.
5.3 Hyperparameters
Method | Top1 F | Top10 F | Uniqueness | Validity | Diversity |
Without MO Selection | 3.830 | 3.791 | 0.999 | 0.816 | 0.842 |
With MO Selection | 4.187 | 4.152 | 0.961 | 0.915 | 0.556 |
To validate the effectiveness of the key components in MOLLM, we conduct a series of ablation studies. In MOLLM, Pareto front selection and F-value selection are applied with equal probability in each iteration. The importance of this design is demonstrated in Table 5, where performance significantly deteriorates when multi-objective selection is removed. Furthermore, if an objective is included in the prompt but is not explicitly considered in MO selection, the performance of MOLLM declines substantially. This highlights the critical role of MO selection in ensuring effective optimization across multiple objectives.
Method | Top1 F | Top10 F | Uniqueness | Validity | Diversity |
GPT-4o direct propose | 3.974 | 3.955 | 0.955 | 0.864 | 0.644 |
MOLLM (Llama3-8B) | 3.988 | 3.900 | 0.986 | 0.482 | 0.749 |
1 offspring each call | 4.068 | 3.980 | 0.969 | 0.942 | 0.575 |
3 offspring each call | 4.208 | 4.114 | 0.970 | 0.831 | 0.592 |
2 offspring each call | 4.276 | 4.245 | 0.949 | 0.900 | 0.529 |
In Table 6, the MOLLM with “2 offspring each LLM call” is used in our official version. Compared to 5000 molecules directly proposed by GPT-4o, MOLLM makes a significant improvement to it, illustrating the effectiveness of our framework. Even with Llama3-8B (Grattafiori et al., 2024) as our backbone, which is much inferior to GPT-4o, its performance is also comparable to other models in Table 1. We make the LLM to generate two offsprings in both crossover and mutation for each LLM call. This design significantly reduces the number of LLM calls needed and achieves better performance, compared to one offspring each call which is used by MOLLEO and three offspring each call.
@@ 6 Conclusion
In this work, we introduce MOLLM, a novel framework that integrates MOO, GA, and LLMs with in-context learning and prompt engineering for molecular design. MOLLM requires no additional training, relying exclusively on LLMs as genetic operators, and achieves SOTA performance in unconstrained molecular optimization. Through rigorous framework design, empirical evaluations, and ablation studies, we demonstrate its effectiveness and efficiency. MOLLM significantly reduces computational costs while outperforming other LLM-based approaches and other SOTA methods. This efficiency is particularly advantageous for practical applications, where molecular property evaluations often involve costly biological and pharmaceutical testing, and LLM inference imposes a high computational overhead. Our results show that MOLLM maintains robust performance across various objective settings and remains superior when optimizing multiple objectives, including less predictable properties such as BBBP. Furthermore, MOLLM is adaptable to different LLM architectures, facilitated by a carefully designed prompt template that fully utilizes LLM knowledge. Future research may focus on enhancing molecular diversity and refining the experience pool mechanism to further improve optimization performance.
Impact Statement
The development of MOLLM introduces a novel approach to multi-objective molecular design by integrating LLMs as genetic operators. This work has the potential to advance computational drug discovery, materials science, and chemical engineering by significantly improving the efficiency and effectiveness of molecular optimization.
From an ethical perspective, MOLLM does not generate molecules directly aimed at harmful applications, such as toxic or hazardous compounds. However, as with any generative model in molecular design, dual-use concerns may arise, necessitating responsible usage and safeguards to ensure ethical deployment. Researchers and practitioners leveraging MOLLM should carefully consider biosecurity implications, regulatory frameworks, and best practices in molecular design.
On a societal level, the framework reduces the reliance on resource-intensive molecular synthesis and experimental testing, potentially accelerating drug discovery and enabling more cost-effective pharmaceutical and material innovations. Additionally, MOLLM’s adaptability to different LLM architectures ensures that future advancements in AI models can further enhance molecular design without requiring retraining or additional computational resources.
Future work should focus on improving molecular diversity, refining the experience pool mechanism, and ensuring ethical guidelines are upheld in real-world applications. This work aligns with the broader goal of advancing machine learning for scientific discovery, contributing to AI-driven molecular design with potential long-term benefits in healthcare, sustainability, and materials innovation.

### v2

@@ 5.2 Experience Pool
Top1 F | Top10 F | AUC-Top10 | Hypervolume | Uniqueness | Validity | Diversity | |
0.0 | 4.276 | 4.245 | 3.992 | 0.867 | 0.949 | 0.900 | 0.529 |
0.3 | 4.354 | 4.260 | 4.004 | 0.881 | 0.879 | 0.905 | 0.509 |
0.5 | 4.413 | 4.388 | 4.116 | 0.905 | 0.872 | 0.908 | 0.441 |
Inspired by LLM planning literature such as ReAct Yao et al. (2022) and ExpeL Zhao et al. (2023), we incorporate an experience pool into our algorithm to enhance molecular optimization. The experience pool consists of two key components: (1) valuable insights and important information kept from past experience, and (2) a summary is constructed from both high-performing and suboptimal molecules generated during the optimization process. These are achieved by summarizing information from the top 10 and bottom 10 molecules in each iteration. The worst 10 molecules are selected using a sliding window approach with a stride of 10. Specifically, if in the previous iteration, the worst molecules were extracted from the bottom 10, the next iteration extracts molecules ranked from the bottom 20 to the bottom 10. This mechanism ensures that the experience pool continuously evolves, integrating knowledge from both current and past iterations. However, incorporating experience into every generation step carries a high risk of convergence to local optima, as it overly exploits the currently discovered search directions. To address this concern, we conduct additional experiments by varying the probability of incorporating experience, as shown in Table 6 and determine to use 0.5. A detailed analysis comparing the initial and final experience in the experience pool is provided in Appendix 7.1. We also visualize the convergence behavior of the top-10 and top-100 molecules generated by baseline models in Figure 2. MOLLM exhibits highly stable training curves, consistently outperforming all other models with a substantial margin over the second-best. Notably, even without the experience pool, MOLLM surpasses MOLLEO, highlighting the effectiveness of its other components.
@@ 6 Conclusion
In this work, we introduce MOLLM, a novel framework that integrates MOO, GA, and LLMs with in-context learning and prompt engineering for molecular design. MOLLM requires no additional training, relying exclusively on LLMs as genetic operators, and achieves SOTA performance in multi-objective molecular optimization and PMO benchmark. Through rigorous framework design, empirical evaluations and analysis, we demonstrate its effectiveness and efficiency. MOLLM significantly reduces computational costs while outperforming other LLM-based approaches and other SOTA methods. This efficiency is particularly advantageous for practical applications, where molecular property evaluations often involve costly biological and pharmaceutical testing, and LLM inference imposes a high computational overhead. Our results show that MOLLM maintains robust performance across various objective settings and remains superior when optimizing many objectives, including less predictable properties such as BBBP. Furthermore, MOLLM is adaptable to various open-source and closed-source LLMs and can be readily extended to other domains with minor modifications to the our code. The source code will be released as open source upon acceptance of the paper to facilitate reproducibility and further research. Due to the limit of pages, we will also provide the ADMET analysis of MOLLM upon acceptance of the paper. However, one limitation of our work is that relying solely on a single LLM to generate new molecules may increase the risk of convergence to local optima. Future research may focus on enhancing the diversity of search directions and improving the exploration efficiency of MOLLM.